OpenStudy (anonymous):

2x4+4x6+6x8.........to n terms..... how to get the 20th term of this sequence

OpenStudy (anonymous):

Un=4n^2+4n U20=4(20)^2+4(20) =1680

OpenStudy (anonymous):

Un=4n(n+1)

OpenStudy (anonymous):

how to reach that equation??

OpenStudy (amistre64):

8 24 48 ... 16 24 8 maybe, then we take the first diagonal and apply some pascal to it with any luck

OpenStudy (anonymous):

When we do sequences(or series) we always need to take the term number into our thinking process. So, naturally, we look at the first term and say 2 times one is 2 so 2n is part of the equation somehow.

OpenStudy (anonymous):

sry, i have a way of doing it, but its not that easy to explain...once you do lots of series and sequences, there tends to be a patter like usually n+1 or n+2 or 2n is involved. Just checks work with my method.

OpenStudy (amistre64):

8*nC0+16*nC1+8*nC2 this is similar to what i recall; prolly needs to be refreshed to be accurate tho

OpenStudy (amistre64):

8*1+16n+8n(n-1)/2 lets see if this is a good recollection lol 8 + 16n +4n^2-4n 4n^2 -12n +8 ?

OpenStudy (amistre64):

when n=0, we get 8 when n=1 we get 0 ... doh!!

OpenStudy (amistre64):

2,4,6 = 2+2(n-1) x 4,6,8 = 4+2(n-1) an = (2+2n-2) * (4+2n-2) a20 = (2+2(20)-2)*(4+2(20)-2) = (2(20))*(2+2(20)) = (40)*(42) = 1680 ?? yay, we match now

OpenStudy (anonymous):

right,,,thanx mates